# Terahertz Landau level spectroscopy of Dirac fermions in millimeter-scale twisted bilayer graphene

### v1

@@ I Introduction
Landau levels form in a magnetic field when electrons undergo cyclotron motion and form states with quantized energy and angular momentum in accordance with the Bohr-Sommerfeld quantization relation. For Dirac fermions, Landau levels are spaced as $\sqrt{B}$ as was demonstrated in scanning tunneling microscopy (STM) probes [40, 41, 42, 43] and infrared spectroscopy [44] of monolayer graphene shortly after it was first isolated [45]. Large area epitaxial multilayer graphene, up to 100 layers thick, with stacking faults that restored the monolayer-like electronic structure later enabled more extensive infrared spectroscopy studies which reconfirmed the $\sqrt{B}$ dependence [46, 47, 48, 49]. In tandem, Landau level spectroscopy of Bernal bilayers was deduced through careful transport measurements [50] and infrared spectroscopy of exfoliated bilayers [51] and “Bernal like” epitaxial multilayer graphene [52] in good agreement with theory [53, 54, 55]. In contrast, twisted bilayer graphene has been inaccessible to free space terahertz probes because of the difficulty in producing samples with a uniform twist angle over a large length scale. Consequently other approaches have been taken to study the low energy electronic structure of TBG. To investigate the gap between Landau levels previous studies have used scanning-tunneling microscopy [43, 56, 57, 58] and quantum oscillation measurements [59, 60, 61] on micrometer-size samples.
Recent progress in ultra-flat substrates and wet-transfer techniques now allows for the preparation of TBG samples that satisfy the sample size requirements of free space terahertz probes [62, 63, 64]. We leverage these techniques to fabricate millimeter-scale twisted bilayer graphene samples over 5 mm $\times$ 5 mm in size with twist angles from $7^{\circ}$ to $20^{\circ}$ with up to a $1^{\circ}$ twist angle inhomogeneity. These samples are compatible with far-field terahertz (THz) spectroscopy, and we conduct time-domain magneto-terahertz polarimetry to resonantly excite inter-Landau level excitations and determine the Fermi velocity. The Fermi velocities we determine are consistent with previous STM and transport studies and thus provide a valuable benchmark of far-field terahertz spectroscopy on a known quantity. This develops terahertz spectroscopy into a tool to study the low energy electronic structure of two-dimensional moiré materials.
@@ II Cyclotron Resonances of Dirac Fermions in Twisted Bilayer Graphene
Figure 1: Electronic structure and doping of large angle twisted bilayer graphene. (a) Spectral function for a $9.43^{\circ}$ twist from a tight-binding model with phenomenological broadening $\eta=0.050$ eV, and (b) calculated density of states with chemical potential consistent with the $9^{\circ}$ experimental sample illustrated. (c) A linear fit to the symmetric part of the measured Hall resistance in the $9^{\circ}$ sample gives the hole carrier density $n_{h}=3.966\times 10^{13}/\mathrm{cm}^{2}$. (d) Zooming in from the gross band structure to the fine structure of Landau levels the levels are separated by a cyclotron frequency $\hbar\omega_{c}$ in the THz range.
@@ III Experimental Study
Figure 2: Illustration of the incident THz beam (green) on the TBG sample and the resulting Faraday rotation. Labeled optical elements include beam splitter (BS), Auston switch (AS), off-axis parabolic mirror (OAP) THz polarizer (P), and delay stage (DS).
Figure 3: Magneto-terahertz measurements of $7^{\circ}$, $9^{\circ}$, $10^{\circ}$ and $20^{\circ}$ twisted bilayer graphene. (a) Real and (b) imaginary parts of the Faraday rotation angle as dependent on magnetic field and light frequency; measured data is the solid lines and fits are the dashed lines. (c) Cyclotron frequency (gap between Landau levels) and (d) scattering rate extracted from fits to the experimental data; a linear fit $\omega_{c}(B)=eB/m^{*}$ of the cyclotron frequencies gives the cyclotron mass $m^{*}$. To extract the effective mass, we only fit data for which $\omega_{c}>\Gamma$ (red points) and exclude data for which $\omega_{c}<\Gamma$ (blue points) so that at least one cyclotron orbit is completed. The scattering rate may vary with magnetic field as the magnetic length scale changes from being much larger than the moiré unit cell to being on the same length scale.
@@
Next, the samples were placed on a 5 mm aperture for THz measurements. The time-domain terahertz setup is outlined in Fig. 2. A 780 nm wavelength femtosecond laser pulse (blue), 85 fs pulses, and 80 MHz repetition rate was split into to arms and sent to two photoconductive terahertz antennas. A delay stage changed the relative arrival time of the optical pulses on the terahertz antennas. The emitter antenna was biased to create current pulses that radiated into free space as broadband terahertz radiation. Four off-axis parabolic mirrors in an 8f focal geometry were used to direct the terahertz pulse to focus on the sample and then to the terahertz detector. When both the terahertz and optical pulse are incident on the detector antenna the produced current is linear in terahertz electric field and this was measured with a lock-in amplifier. Terahertz measurements were made in a Lakeshore cryostat at 1.7 K with a static magnetic field perpendicular to the surface of the sample.
The frequency-dependent Faraday rotation angle was obtained by sending a THz pulse polarized in $\hat{x}$ to the sample and then separately measuring transmitted THz electric fields $E_{x}$ and $E_{y}$. Both electric fields are measured in the time domain and then Fourier transformed to obtain the spectrum. The spectrum was measured from 0.3 to 2.5 THz and was frequency limited by the spectrum of the initial THz pulse. The Faraday spectrum, $\theta_{F}(\omega)$, is then defined as
@@
The Faraday spectrum is a complex quantity. The real component is a polarization rotation of the THz pulse, and the imaginary component describes the ellipticity of the transmitted pulse. The THz detector directly measures the THz electric field not the intensity $\propto E^{2}(\omega)$), so the real and imaginary Faraday angle was directly obtained without using Kramers-Kronig relations.
Our THz setup required three THz polarizers to measure $\theta_{F}$. One was placed before the sample to set the incident polarization. A second polarizer was placed on a motorized stage after the sample. This polarizer discriminates between the $\hat{x}$ and $\hat{y}$ components of the transmitted terahertz pulse. The third polarizer was set at $45^{\circ}$ right before the detector antenna. The terahertz antennas had a polarization preference, so the third polarizer made the detector equally sensitive to the $\hat{x}$ and $\hat{y}$ components [66, 67].
@@ IV.2 Cyclotron Mass
The complex terahertz (THz) Faraday rotation, measured as illustrated in Fig. 3, gives the cyclotron resonance which can be fitted to obtain the cyclotron mass. To analyze the Faraday rotation we follow the approach in Refs. [68, 69, 70] where the complex Faraday rotation is defined as
@@
where $n_{s}$ is the substrate refractive index, $Z_{0}=4\pi\alpha\hbar/e^{2}\approx 376.73\ \Omega$ is the impedance of free space, and $G_{\pm}$ is the conductance under right/left-handed circularly polarized light. We take the conductance to be composed of a Drude-type term shifted by the cyclotron resonance $\omega_{c}$ (gap between Landau levels) and a lattice polarizability $\epsilon_{\infty}$ from absorptions above the measured spectral range
@@
where the term in parentheses is dimensionless, $d$ is the thickness of the bilayer, $\omega_{p}$ is the Drude plasma frequency, and $\Gamma=1/\tau$ is the scattering rate in terms of $\tau$ the time between scattering events.
The measured Faraday rotations are the solid curves in Fig. 3(a-b) and their fits obtained from Eq. 8 are the dashed lines. The cyclotron resonance appears as an inflection point in the real component of the Faraday spectrum and a minimum in the imaginary component. Fitting the cyclotron resonance as a function of magnetic field then gives the effective mass as $m^{*}=eB/\omega_{c}$ as illustrated in Fig. 3(c). Fig. 3(d) shows the scattering rate as dependent on field obtained from the fits. Here we limit ourselves to analyzing data at fields where $\omega_{c}\tau>1$ (red), so that at least a single cyclotron orbit is completed. Landau level transitions with $\omega_{c}\tau<1$ (blue) are not fitted.
@@ IV.3 Fermi Velocity
Combining these data together we can then obtain the Fermi velocity. For TBG we plot the velocities in Fig. 4 and see that all the samples are consistent with the relatively constant Fermi velocity for $\theta\gtrsim 5^{\circ}$ predicted by theory and are in the range of Fermi velocities previously determined using STM and transport probes of small-scale samples. Perhaps somewhat surprisingly the Fermi velocities we determine for all samples are in the expected range despite significant $p$-doping; in particular the chemical potential of the $7^{\circ}$ sample lies near the van Hove singularity so we might expect a breakdown of Dirac physics for this sample. For extended data see Appendix C. Our results demonstrate a new experimental approach to probe the electronic structure of two-dimensional materials and we find that Dirac-like physics persists well away from charge neutrality in twisted bilayer graphene.
Figure 4: Fermi velocities determined by magneto-optical spectroscopy (red), previous STM studies (gray), and quantum oscillations studies (yellow). Lower and upper dashed lines are the theoretical fit from Eq. 4 with $v_{F}^{\mathrm{mono}}=1.0$ and $1.4\times 10^{6}$ m/s respectively and $t_{\perp}=0.1$ eV. Variation in monolayer Fermi velocity can originate from the level of doping [71, 72].
@@ V Discussion
Two dimensional materials have emerged as compelling platforms to realize correlated electronic physics. Until recently this physics has been limited to short length scales of $\sim 10\ \mu\mathrm{m}$, but with advances in materials synthesis, twisted bilayer graphene with a uniform twist angle on the length scale of $\sim 1000\ \mu\mathrm{m}=1\ \mathrm{mm}$ is now available. Leveraging these techniques we interfaced millimeter length scale twisted bilayer graphene with terahertz magneto-optics and conducted complex Faraday rotation and transport measurements to reveal the electronic structure of these materials. We measure the gap between Landau levels and find Fermi velocities that are consistent with previous scanning tunneling microscopy and transport studies thus validating our far-field approach on large samples. This suggests that these macroscopic samples exhibit good coherence over long length scales, a necessary prerequisite to realize reproducible macroscopic quantum materials. Extending bilayer materials from microscopic to mesoscopic and macroscopic length scales enhances scalability for advancing quantum technologies and may enable the realization of other exotic phenomena such as weak localization and universal conductance fluctuations  [^1] D. Mallick, A.H. Chen, S. Talkington, B. F. Mead, S. J. Yang, C. J. Kim, E. J. Mele, L. Wu, and M. Brahlek, Weak localization in twisted bilayer graphene (in prep.). Looking ahead, terahertz techniques are a promising direction to probe collective excitations in two-dimensional materials such as twisted bilayer graphene and transition metal dichalcogenides.
@@ VI Acknowledgments
The THz measurements were mainly supported by L.W.’s startup package at Penn. B.M. was partially sponsored by the Army Research Office under Grant Number W911NF-20-2-0166 and W911NF-25-2-0016. S.T. acknowledges support from the NSF under Grant No. DGE-1845298. A.H.C., D.M., and M.B. were supported by the U. S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division, and the National Quantum Information Science Research Centers, Quantum Science Center. X.H. acknowledges the partial support by Air Force Office of Scientific Research under award no. FA955022-1-0410 and the NSF EPM program under grant no. DMR-2213891. S.J.Y. and C.J.K. acknowledge support from the Institute for Basic Science (IBS), Korea, under Project Code IBS-R034-D1. E.J.M. was funded by the Department of Energy under grant DE-FG02-84ER45118. L.W. acknowledges support from the Sloan Foundation under the award FG-2025-25036.

### v2

@@ I Introduction
Landau levels form in a magnetic field when electrons undergo cyclotron motion and form states with quantized energy and angular momentum in accordance with the Bohr-Sommerfeld quantization relation. For Dirac fermions, Landau levels are spaced as $\sqrt{B}$ as was demonstrated in scanning tunneling microscopy (STM) probes [40, 41, 42, 43] and infrared spectroscopy [44] of monolayer graphene shortly after it was first isolated [45]. Large area epitaxial multilayer graphene, up to 100 layers thick, later reconfirmed the $\sqrt{B}$ dependence as stacking faults restored the monolayer-like electronic structure [46, 47, 48, 49]. In tandem, Landau level spectroscopy of Bernal bilayers was deduced through careful transport measurements [50] and infrared spectroscopy of exfoliated bilayers [51] and “Bernal like” epitaxial multilayer graphene [52] in good agreement with theory [53, 54, 55]. In contrast, twisted bilayer graphene has been inaccessible to free space terahertz probes because of the difficulty in producing samples with a uniform twist angle over a large length scale. Consequently other approaches have been taken to study the low energy electronic structure of TBG. To investigate the gap between Landau levels previous studies have used scanning-tunneling microscopy [43, 56, 57, 58] and quantum oscillation measurements [59, 60, 61] on micrometer-size samples.
Recent progress in ultra-flat substrates and wet-transfer techniques now allows for the preparation of TBG samples that satisfy the sample size requirements of free space terahertz probes [62, 63, 64]. We leverage these techniques to fabricate millimeter-scale twisted bilayer graphene samples up to 5 mm $\times$ 5 mm in size with twist angles from $7^{\circ}$ to $20^{\circ}$ with up to a $1^{\circ}$ twist angle inhomogeneity. These samples are compatible with far-field terahertz spectroscopy, and we conduct time-domain magneto-terahertz polarimetry to resonantly excite inter-Landau level excitations and determine the Fermi velocity. The Fermi velocities we determine are consistent with previous STM and transport studies and thus provide a valuable benchmark of far-field terahertz spectroscopy on a known quantity. This develops terahertz spectroscopy into a tool to study the low energy electronic structure of two-dimensional moiré materials.
@@ II Cyclotron Resonances of Dirac Fermions in Twisted Bilayer Graphene
Figure 1: Electronic structure and doping of large angle twisted bilayer graphene. (a) Spectral function for a $9.43^{\circ}$ twist from a tight-binding model with phenomenological broadening $\eta=0.050$ eV, and (b) calculated density of states with chemical potential consistent with the $9^{\circ}$ experimental sample illustrated. (c) A linear fit to the asymmetric part of the measured Hall resistance in the $9^{\circ}$ sample gives the hole carrier density $n_{h}=3.966\times 10^{13}/\mathrm{cm}^{2}$. (d) Zooming in from the gross band structure to the fine structure of Landau levels the levels are separated by a cyclotron frequency $\hbar\omega_{c}$ in the terahertz range.
@@ III Experimental Study
Figure 2: Illustration of the incident terahertz beam (green) on the TBG sample and the resulting Faraday rotation. Labeled optical elements include beam splitter (BS), Auston switch (AS), off-axis parabolic mirror (OAP) terahertz polarizer (P), and delay stage (DS).
Figure 3: Magneto-terahertz measurements of $7^{\circ}$, $9^{\circ}$, $10^{\circ}$ and $20^{\circ}$ twisted bilayer graphene. (a) Real and (b) imaginary parts of the Faraday rotation angle as dependent on magnetic field and light frequency; measured data is the solid lines and fits are the dashed lines. The samples are measured by terahertz pulses where the entire frequency range simultaneously probes the sample, however, the intensity of the pulse is much smaller at the extreme ends of the spectral range which results in noise at the upper end of the frequency range in the case $9^{\circ}$ sample. (c) Cyclotron frequency (gap between Landau levels) and (d) scattering rate extracted from fits to the experimental data; a fit to the formula $\omega_{c}=(e/m^{*})\sqrt{B_{i}^{2}+B^{2}}$ gives the cyclotron mass $m^{*}$ and a length-scale $L_{i}=\sqrt{\hbar/eB_{i}}$ corresponding to disorder in the samples.
@@
Next, the samples were placed on a 5 mm aperture for terahertz measurements. The time-domain terahertz setup is outlined in Fig. 2. A 780 nm wavelength femtosecond laser pulse (blue), 85 fs pulses, and 80 MHz repetition rate was split into to arms and sent to two photoconductive terahertz antennas. A delay stage changed the relative arrival time of the optical pulses on the terahertz antennas. The emitter antenna was biased to create current pulses that radiated into free space as broadband terahertz radiation. Four off-axis parabolic mirrors in an 8f focal geometry were used to direct the terahertz pulse to focus on the sample and then to the terahertz detector. When both the terahertz and optical pulse are incident on the detector antenna the produced current is linear in terahertz electric field and this was measured with a lock-in amplifier. Terahertz measurements were made in a Lakeshore cryostat at 1.7 K with a static magnetic field perpendicular to the surface of the sample.
The frequency-dependent Faraday rotation angle was obtained by sending a terahertz pulse polarized in $\hat{x}$ to the sample and then separately measuring transmitted terahertz electric fields $E_{x}$ and $E_{y}$. Both electric fields are measured in the time domain and then Fourier transformed to obtain the spectrum. The spectrum was measured from 0.3 to 2.5 THz and was frequency limited by the spectrum of the initial terahertz pulse. The Faraday spectrum, $\theta_{F}(\omega)$, is then defined as
@@
The Faraday spectrum is a complex quantity. The real component is a polarization rotation of the terahertz pulse, and the imaginary component describes the ellipticity of the transmitted pulse. The terahertz detector directly measures the terahertz electric field, not the intensity which is proportional to $E^{2}(\omega)$, so the real and imaginary Faraday angle was directly obtained without using Kramers-Kronig relations.
Our terahertz setup required three terahertz polarizers to measure $\theta_{F}$. One was placed before the sample to set the incident polarization. A second polarizer was placed on a motorized stage after the sample. This polarizer discriminates between the $\hat{x}$ and $\hat{y}$ components of the transmitted terahertz pulse. The third polarizer was set at $45^{\circ}$ right before the detector antenna. The terahertz antennas had a polarization preference, so the third polarizer made the detector equally sensitive to the $\hat{x}$ and $\hat{y}$ components [66, 67].
@@ IV.2 Cyclotron Mass
The complex terahertz Faraday rotation, measured as illustrated in Fig. 3, gives the cyclotron resonance which can be fitted to obtain the cyclotron mass. To analyze the Faraday rotation we follow the approach in Refs. [68, 69, 70] where the complex Faraday rotation is defined as
@@
where $n_{s}$ is the substrate refractive index, $Z_{0}=4\pi\alpha\hbar/e^{2}\approx 376.73\ \Omega$ is the impedance of free space, and $G_{\pm}$ is the conductance under right/left-handed circularly polarized light. We take the conductance to be composed of a Drude-type term shifted by the cyclotron resonance and a lattice polarizability from absorptions above the measured spectral range
@@
where the term in parentheses is dimensionless, $d$ is the thickness of the bilayer, $\omega_{p}$ is the Drude plasma frequency, $\omega_{c}$ is the cyclotron resonance (gap between Landau levels), $\Gamma=2\pi/\tau$ is the scattering rate, $\tau$ is the average time between scattering events, and $\varepsilon_{\infty}$ is the lattice polarizability.
The measured Faraday rotations are the solid curves in Fig. 3(a-b) and their fits obtained from Eq. 8 are the dashed lines. The cyclotron resonance appears as an inflection point in the real component of the Faraday spectrum and a minimum in the imaginary component. The 1 T data for the $20^{\circ}$ sample was omitted because the measured Faraday spectrum had oscillations that could not be fit with Eq. 10. Now, due to disorder effects, $\omega_{c}=eB/m^{*}$ is inaccurate at low fields. A fit to $\omega_{c}(B)$ that captures minimal features of disorder is the formula $\omega_{c}=\frac{e}{m^{*}}\sqrt{B_{i}^{2}+B^{2}}$ which has a non-zero asymptote corresponding to disorder scattering processes [71]. Fitting the cyclotron resonance as a function of magnetic field, Fig. 3(c), gives the effective mass $m^{*}$ and impurity scattering length $L_{i}=\sqrt{\hbar/eB_{i}}$. We find $L_{i}$ of 13-21 nm for these samples, which is consistent with the length-scale for inter-valley scattering [74]. Fig. 3(d) shows the scattering rate vs. magnetic field obtained from the fits.
@@ IV.3 Fermi Velocity
We obtain the Fermi velocity by combining the transport and terahertz results. For TBG we plot the velocities in Fig. 4 and see that all the samples are consistent with the relatively constant Fermi velocity for $\theta\gtrsim 5^{\circ}$ predicted by theory and are in the range of Fermi velocities previously determined using STM and transport probes of small-scale samples. Perhaps somewhat surprisingly the Fermi velocities we determine for all samples are in the expected range despite significant $p$-doping; in particular the chemical potential of the $7^{\circ}$ sample lies near the van Hove singularity so we might expect a breakdown of Dirac physics for this sample. Now if twist angle was the only difference between samples we would expect the $9^{\circ}$ and $10^{\circ}$ to be nearly indistinguishable, however our samples have different chemical potentials, mobilities, and disorder profiles leading to the observed Fermi velocities. For extended data see Appendix C. Our results demonstrate a new experimental approach to probe the electronic structure of two-dimensional materials and we find that Dirac-like physics persists well away from charge neutrality in twisted bilayer graphene.
Figure 4: Fermi velocities determined by magneto-optical spectroscopy (red), previous STM studies (gray), and quantum oscillations studies (yellow). Lower and upper dashed lines are the theoretical fit from Eq. 4 with $v_{F}^{\mathrm{mono}}=1.0$ and $1.2\times 10^{6}$ m/s respectively and $t_{\perp}=0.1$ eV. Variation in monolayer Fermi velocity can originate from the level of doping [72, 73].
@@ V Discussion
Two dimensional materials have emerged as compelling platforms to realize correlated electronic physics. Until recently this physics has been limited to short length scales of $\sim 10\ \mu\mathrm{m}$, but with advances in materials synthesis, twisted bilayer graphene with a uniform twist angle on the length scale of $\sim 1000\ \mu\mathrm{m}=1\ \mathrm{mm}$ is now available. Leveraging these techniques we interfaced millimeter length scale twisted bilayer graphene with terahertz magneto-optics and conducted complex Faraday rotation and transport measurements to reveal the electronic structure of these materials. We measure the gap between Landau levels and find Fermi velocities that are consistent with previous scanning tunneling microscopy and transport studies thus validating our far-field approach on large samples. This suggests that these macroscopic samples exhibit good coherence over long length scales, a necessary prerequisite to realize reproducible macroscopic quantum materials. Extending bilayer materials from microscopic to mesoscopic and macroscopic length scales enhances scalability for advancing quantum technologies and may enable the realization of other exotic phenomena such as weak localization and universal conductance fluctuations  [^1] S. Talkington, D. Mallick, A.H. Chen, B. F. Mead, S. J. Yang, C. J. Kim, S. Adam, L. Wu, M. Brahlek, and E. J. Mele, Weak localization in twisted bilayer graphene (in prep.). Looking ahead, terahertz techniques are a promising direction to probe collective excitations in two-dimensional materials such as twisted bilayer graphene and transition metal dichalcogenides.
@@ VI Acknowledgments
The terahertz measurements were mainly supported by L.W.’s startup package at Penn. B.M. was partially sponsored by the Army Research Office under Grant Number W911NF-20-2-0166 and W911NF-25-2-0016. S.T. acknowledges support from the NSF under Grant No. DGE-1845298. A.H.C., D.M., and M.B. were supported by the U. S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division, and the National Quantum Information Science Research Centers, Quantum Science Center. X.H. acknowledges the partial support by Air Force Office of Scientific Research under award no. FA955022-1-0410 and the NSF EPM program under grant no. DMR-2213891. S.J.Y. and C.J.K. acknowledge support from the Institute for Basic Science (IBS), Korea, under Project Code IBS-R034-D1. E.J.M. was funded by the Department of Energy under grant DE-FG02-84ER45118. L.W. acknowledges support from the Sloan Foundation under the award FG-2025-25036.